OpenStudy (anonymous):

Which equations could model this problem? A 5 lb bag of potatoes costs $3. What is the unit price p per pound? Choose all answers that are correct. A. B. C. D.

OpenStudy (anonymous):

A.3/5=P B.3P=5 C.5/3=P D.5P=3

OpenStudy (igreen):

To find out how much 1 pound costs divide 5 and 3. 3 / 5 = 0.60 So 1 pound of potatoes will be $0.60

OpenStudy (igreen):

In other words your answer is A.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @almightyemery

OpenStudy (anonymous):

can u help me wit more

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

its only A

OpenStudy (igreen):

Yes, the answer is A.

OpenStudy (igreen):

I can help with more, just close this question and open another one.

OpenStudy (anonymous):

A is correct. However, you might want to look at answer D too.